Write an equation in standard form for the circle that statisfies these conditions
center(0,7) tangent to x-axis Answer by lwsshak3(11628) 
 Write an equation in standard form for the circle that statisfies these conditions
center(0,7) tangent to x-axis
**
If you make a rough sketch of given points, you will see that the radius=7
Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius.
Equation of given circle
x^2+(y-7)^2=7^2
x^2+(y-7)^2=49